OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 2x + 4.

OpenStudy (ivettef365):

convert the equation as y = mx + b

OpenStudy (ivettef365):

where m = slope

OpenStudy (ivettef365):

just divide all digits by 8

OpenStudy (anonymous):

- 8y = 2x+4 - y = (2x+4)/8 - y = 2(x+2)/8 - simplifies to (x+2)/4 - which can be rewritten as y = (x/4)+(2/4) - which is also equivalent to y = 0.25x + 0.5 if y = mx+b is your standard equation, m represents slope thus 0.25 (or 1/4) is your slope value